# Capivara: A Spectral-based Segmentation Method for IFU Data Cubes

### v1

@@ 4.2 Emission lines classification
After subtracting the stellar spectra derived in section 4.1, we measured the flux of the four main emission lines (H, H, [Nii]6583Å and [Oiii]5007Å) required to build the Baldwin-Phillips-Terlevich diagram (BPT; Baldwin et al., 1981; Veilleux & Osterbrock, 1987; Rola et al., 1997; Kewley et al., 2001; Kauffmann et al., 2003; Stasińska, 2007; Schawinski et al., 2010).This analysis was conducted by fitting a Gaussian function to each emission line using the ifscube code (Ruschel-Dutra et al., 2021). ifscube fits multiple Gaussian or Gauss-Hermite profiles, with or without constraints, and supports pixel-by-pixel uncertainties, weights, flags, and pseudo-continuum fitting. The best fit is determined by minimizing the , with free parameters being the amplitude, radial velocity, and velocity dispersion of each component. Our sample galaxies exhibited no signs of multiple components per emission line, and a single narrow Gaussian was sufficient. Given the similar ionization potentials of H, H and [NII] have similar ionization potentials, their kinematic properties were fixed, while [OIII] was allowed independent radial velocity and velocity dispersion. Fig 6, provides an example fit for galaxy 10224-6104 at (100, 00).
The H, H, [Nii], and [Oiii] maps are displayed in Figures 7 and 8. From left to right, the panels depict the fluxes for [O iii], H, [N ii], and H. The top row presents the fluxes measured from the original data cube, while the bottom row shows the measurements after running capivara. Similar to the stellar continuum, our method effectively identifies the main emission-line regions in each object. However, the improvement in S/N is much more pronounced in the emission lines. In many regions of the original datacube, these features are not properly measured due to not meeting the S/N threshold. This issue is resolved in the decomposed datacube, where our model not only enhances the S/N and correctly groups the regions but also ensures more consistent and accurate measurements of these features, without merging regions that are not physically meaningful from a spectral perspective.
@@ 4.3 Continuum and emission line spectral classification
Based on the dominant stellar populations in each object, we divided our sample into six classes: AGN, star-forming (SF), post-starburst (PS), intermediate-age (IA), mixed (MIX), and quenched (QUENCH). The AGN class includes objects dominated by a power-law-like spectrum, while the star-forming class includes objects dominated by simple stellar populations (SSPs) younger than 100 Myr. The post-starburst class consists of objects dominated by SSPs between 100 and 700 Myr, while the intermediate-age class consists of SSPs between 700 Myr and 2 Gyr. The mixed class includes objects where SSPs older than 2 Gyr contribute less than 40% of the luminosity, and the quenched class includes objects where SSPs older than 2 Gyr contribute more than 40% of the luminosity.
@@
The BPT diagram classifies galaxies into five distinct categories based on their emission line fluxes: inactive (INACT), star-forming (SF), composite or transition (TRANSIT), low-ionization nuclear emission-line region (LINER), and Seyfert galaxies. It utilizes the flux ratios of H, [OIII] 5007, H, and [NII] 6583, plotting galaxies in a parameter space defined by on the x-axis and on the y-axis. The thresholds between these line ratios, as defined by Kewley et al. (2001) and Kauffmann et al. (2003), are used to differentiate between these classes. Inactive galaxies are identified by [Oiii] emission below 3- of the continuum signal-to-noise ratio (S/N). Star-forming galaxies, primarily influenced by photoionization from young, hot stars, exhibit low values of both [Nii]/H and [Oiii]/H, occupying the left wing locus of the BPT diagram. In contrast, AGN-dominated galaxies, including LINERs and Seyferts, lie on the opposite side. LINERs display high [Nii]/H but low [Oiii]/H ratios, with their ionization source being uncertain, possibly due to nuclear activity or evolved stellar populations. Seyferts show high values in both [Nii]/H and [Oiii]/H ratios. Finally, the transition class falls in the zone between star-forming and AGN objects, demarcated by the theoretical extreme starburst line proposed by Kewley et al. (2001) and the empirical starburst line by Kauffmann et al. (2003).
Figure 9 show the classification maps for our five galaxies. The first two panels on the left in each row display the maps for the continuum classes, while the two panels on the right present the classification based on emission lines. In both cases, the original datacube is shown first, followed by the decomposed datacube. Although a more detailed analysis of subtle differences is beyond the scope of this initial paper, we note an overall agreement between the segmented regions detected by capivara and the predefined classes. Future studies will aim to refine this approach, tailoring it to more domain-specific cases.
@@ 5 Conclusions
Our method, however, takes a different approach. Primarily designed to decompose data into regions with distinct physical properties, it naturally enhances the S/N as a secondary benefit. In contrast to Voronoi binning, which focuses on data quality by merging regions that may have different physical properties, our approach ensures that only regions with similar characteristics are grouped. In our current analysis, decomposing the data cube into 20 distinct regions results in a five-fold improvement in S/N compared to individual spectra.666This estimate was derived from the continuum between 5500 and 5600 Å, chosen for its minimal stellar absorption and emission lines.
@@
In a demonstration using five selected MaNGA galaxies, capivara identified and grouped regions with similar physical properties. These groupings were validated against standard methods for stellar continuum and gas property analysis. By providing both decomposed data cubes and integrated one-dimensional spectra for each region, capivara facilitates a more detailed investigation of stellar and ionized gas properties.
@@ Acknowledgements
LGDH acknowledges support by National Key R&D Program of China No.2022YFF0503402, and National Natural Science Foundation of China (NSFC) project number E345251001.

### v2

@@ 4.2 Emission lines classification
After subtracting the stellar spectra derived in section 4.1, we measured the flux of the four main emission lines (H, H, [Nii]6583Å and [Oiii]5007Å) required to build the Baldwin-Phillips-Terlevich diagram (BPT; Baldwin et al., 1981; Veilleux & Osterbrock, 1987; Rola et al., 1997; Kewley et al., 2001; Kauffmann et al., 2003; Stasińska, 2009; Schawinski et al., 2010). This analysis was conducted by fitting a Gaussian function to each emission line using the ifscube code (Ruschel-Dutra et al., 2021). ifscube fits multiple Gaussian or Gauss-Hermite profiles, with or without constraints, and supports pixel-by-pixel uncertainties, weights, flags, and pseudo-continuum fitting. The best fit is determined by minimizing the , the free parameters being the amplitude, radial velocity, and velocity dispersion of each component. Our sample galaxies exhibited no signs of multiple components per emission line, and a single narrow Gaussian was sufficient. Given that similar ionization potentials of H, H and [NII] have similar ionization potentials, their kinematic properties were fixed, while [OIII] was allowed independent radial velocity and velocity dispersion. Fig 8, provides an example fit for galaxy 10224-6104 at (100, 00).
The H, H, [Nii], and [Oiii] maps are displayed in Figures 9 and 10. From left to right, the panels depict the fluxes for [O iii], H, [N ii], and H. The top row presents the fluxes measured from the original data cube, while the bottom row shows the measurements after running capivara. Similar to the stellar continuum, our method effectively identifies the main emission-line regions in each object. However, the improvement in S/N is much more pronounced in the emission lines. In many regions of the original datacube, these features are not properly measured due to not meeting the S/N threshold. This issue is resolved in the decomposed datacube, where our model not only enhances the S/N and correctly groups the regions but also ensures more consistent and accurate measurements of these features, without merging regions that are not physically meaningful from a spectral perspective.
@@ 4.3 Continuum and emission line spectral classification
We divided our sample into five classes according to the dominant stellar populations in each object, presented here in the order of appearance: AGN, star-forming (SF), post-starburst (PS), mixed (MIX), and quenched (QUENCH). The AGN class includes objects dominated by a FC component. The SF class consists of objects where simple stellar populations (SSPs) younger than 100 Myr dominate the emission. The PS class is defined by objects in which SSPs between 100 and 2000 Myr dominate. In the MIX class, SSPs older than 2 Gyr contribute less than 40% of the luminosity, while in the QUENCH class, SSPs older than 2 Gyr contribute more than 40% of the luminosity.
@@
The BPT diagram classifies galaxies into five distinct categories based on their emission-line fluxes: inactive (INACT), star-forming (SF), composite or transition (TRANSIT), low-ionization nuclear emission-line region (LINER), and Seyfert galaxies. Here, we utilize the BPT diagram based on the flux ratios of H, [OIII] 5007, H, and [NII] 6583, plotting galaxies in a parameter space defined by on the x-axis and on the y-axis. The thresholds between these line ratios, as defined by Kewley et al. (2001) and Kauffmann et al. (2003), are used to differentiate between these classes. Inactive galaxies are identified by [Oiii] emission below 3- of the continuum signal-to-noise ratio (S/N). Star-forming galaxies, primarily influenced by photoionization from massive young hot stars, exhibit low values of both [Nii]/H and [Oiii]/H, occupying the left wing locus of the BPT diagram. In contrast, AGN-dominated galaxies, including LINERs and Seyferts, lie on the opposite side. LINERs display high [Nii]/H but low [Oiii]/H ratios, with their ionization source being uncertain, possibly due to nuclear activity or evolved stellar populations. Seyferts show high values in both [Nii]/H and [Oiii]/H ratios. Finally, the transition class falls in the zone between star-forming and AGN objects, demarcated by the theoretical extreme starburst line proposed by Kewley et al. (2001) and the empirical starburst line by Kauffmann et al. (2003).
Figure 11 shows the classification maps for our five galaxies. The first two panels on the left in each row display the maps for the continuum classes, while the two on the right present the classifications based on emission lines. In both cases, the original datacube is shown first, followed by the segmented version. Before the capivara segmentation, many galaxies exhibit low S/N regions where classifications fluctuate across adjacent spaxels, often within areas smaller than the PSF. These fluctuations are not physically meaningful, as coherent structures should span spatial scales larger than the PSF. While our method does not explicitly enforce spatial coherence, it preserves spectral similarity across regions, which naturally leads to more spatially coherent segmentations. This, in turn, mitigates the likelihood of merging physically distinct regions. Nevertheless, in some cases, adjacent regions with subtle spectral differences may still be grouped together, leading to the dilution or suppression of less dominant classes. This limitation can be alleviated by increasing the number of classes, though the optimal choice is likely galaxy-dependent and best addressed in dedicated applications. Even so, the segmented regions often align with physically distinct components—whether driven by gas excitation, stellar populations, or a combination of both.
@@ 5 Conclusions
Our method, however, takes a different approach. Primarily designed to segment data into regions with distinct physical properties, it naturally enhances the S/N as a secondary benefit. In contrast to Voronoi binning—which focuses on data quality by merging regions that may have different physical characteristics—our approach ensures that only regions with similar properties are grouped. In our current analysis, segmenting the data cube into 20 distinct regions results in a five-fold improvement in S/N compared to individual spectra.888This estimate was derived from the continuum between 5500 and 5600 Å, chosen for its minimal stellar absorption and emission lines.
@@
As a caveat, the quality of the segmentation produced by capivara ultimately depends on the spectral features present in the data. This is not a limitation of the method itself, but rather a reflection of the information content available to the algorithm. Users should consider pre-processing the spectra for specific needs, emphasizing or isolating regions of interest—such as prominent emission or absorption features—or adopting more physically motivated distance metrics, depending on the scientific goal. Another current limitation of our package is the lack of correction for internal kinematics, such as velocity shifts across the galaxy. While this does not appear to significantly impact the results for the systems analyzed—since spectral variations from ionization or stellar populations tend to dominate—it may become more relevant in galaxies with strong velocity gradients. Future incarnations of the code may address this effect through case-by-case implementations that incorporate kinematic effects into the clustering process.
In a demonstration using five selected MaNGA galaxies, capivara identified and grouped regions with similar physical properties. These groupings were validated against standard methods for stellar continuum and gas property analysis. By providing both segmented data cubes and integrated one-dimensional spectra for each region, capivara facilitates a more detailed investigation of stellar and ionized gas properties.
@@ Acknowledgements
R. S. de Souza acknowledges support from the São Paulo Research Foundation (FAPESP), project 2024/05315-4. L.G.D.H. acknowledges support from the National Key R&D Program of China (2022YFF0503402) and the National Natural Science Foundation of China (NSFC, E345251001). A.C.S. acknowledges funding from the Brazilian National Council for Scientific and Technological Development (CNPq, grants 314301/2021-6 and 445231/2024-6) and the Rio Grande do Sul State Research Foundation (FAPERGS, 24/2551-0001548-5). P.C. acknowledges support from CNPq (310555/2021-3) and FAPESP (2024/05315-4). A.C.S. and R.R. also acknowledge support from the Coordination for the Improvement of Higher Education Personnel (CAPES, project 0001). R.R. further acknowledges support from CNPq (404238/2021-1, 445231/2024-6) and FAPERGS (24/2551-0001282-6), as well as from CAPES (88881.109987/2025-01).